What is the area and perimeter of a square with a diagnol of 10 inche?
Answer by mananth(16946)

diagonal of square = 10 in.
let the side be x in.
The diagonal and the sides of square form a right angle.
Apply pythagoras theorem
x^2 +x^2 = 10^2
2x^2 =100
x^2=50
x= 7.07 in
Area = 50 sq. in
perimeter = 4*7.07
=28.28 inches